In a company, 55% of the workers are women. If 2,475 men work for the company, how many workers are there in all? Use pencil and paper. Show two different ways that you can solve this problem.
There

Answer by stanbon(75887)

In a company, 55% of the workers are women. If 2,475 men work for the company, how many workers are there in all?
Let "x" be the number of workers.
Equation:
# of women + # of men = x
0.55x + 2475 = x
0.45x = 2475
x = 2475/0.45
x = 5500
----------------------
If 55% are women, 45% must be men.
Equation:
9.45x = 2475
x = 2475/0.45
x = 5500
====================
Cheers,
Stan H.
------------
Use pencil and paper. Show two different ways that you can solve this problem.